CHIRPED-PULSE FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF MET A-CHLOROBENZALDEHYDE

\underline{SEAN T. ARNOLD}, JESSICA A. GARRETT, and GORDON G. BROWN, Department of Science and Mathematics, Coker College, 300 E College Ave., Hartsville, SC 29550..

The pure rotational spectrum of meta-chlorobenzaldehyde (m-ClBA) has been measured from 8 - 18.5 GHz by chirpedpulse Fourier transform microwave (CP-FTMW) spectroscopy. The spectrum has been analyzed to discover the presence of two conformations of m-ClBA in the free jet expansion. For each conformation the rotational constants, the centrifugal distortion constants, and the nuclear quadrupole coupling constants have been found for both the ^{35}Cl and the ^{37}Cl isotopologue. The rotational constants and the nuclear quadrupole coupling constants have been compared to ab initio calculations performed using the Gaussian 03W software package.